The lifetime X in hours of a certain electrical component has the pdf

If a random sample of 36 is taken from these components,

find P(3\overline{X} + 100 < 111)
Answer #1

Here X has exponential distribution with parameter 1/3 So

mu=E(X)=3,sigma=3

Since sample size n=36 is greater than 30 so according to CLT sampling distribution of sample mean will be approximately normal with mean and SD as follows:

mu_{ar{x}}=mu=3,sigma_{ar{x}}=rac{sigma}{sqrt{n}}=rac{3}{sqrt{36}}=0.5

Now,

P(3ar{X}+100<111)=P(3ar{X}<11)=P(ar{X}<11/3)

The z-score for T 11/3 is

z=rac{ar{x}-mu_{ar{x}}}{sigma_{ar{x}}}=rac{11/3-3}{0.5}=1.33

Therefore required probability is

P(3ar{X}+100<111)=P(z<1.33)=0.9082

Answer: 0.9082